OpenStudy (diyadiya):

if f(1)=-2 ,f'(1)=1, Find the derivative of logf(e^x) wrt x at x=0

OpenStudy (anonymous):

assuming log = ln(e^x)

OpenStudy (anonymous):

\[\frac{d}{dx}\log(f(e^x))=\frac{f'(e^x)\times e^x}{f(e^x)}\] by the chain rule.

OpenStudy (anonymous):

so at \[x=0\] you get \[\frac{d}{dx}\log(f(e^0))=\frac{f'(e^0)\times e^0}{f(e^0)}=\frac{f'(1)\times 1}{f(1)}\] so now plug in the values you know

OpenStudy (diyadiya):

oh ok Thankyou :)